# SCA: Highly Efficient Semantic-Consistent Unrestricted Adversarial Attack

### v2

@@ Abstract
Unrestricted adversarial attacks typically manipulate the semantic content of an image (e.g., color or texture) to create adversarial examples that are both effective and photorealistic. Recent works have utilized the diffusion inversion process to map images into a latent space, where high-level semantics are manipulated by introducing perturbations. However, they often results in substantial semantic distortions in the denoised output and suffers from low efficiency. In this study, we propose a novel framework called Semantic-Consistent Unrestricted Adversarial Attacks (SCA), which employs an inversion method to extract edit-friendly noise maps and utilizes Multimodal Large Language Model (MLLM) to provide semantic guidance throughout the process. Under the condition of rich semantic information provided by MLLM, we perform the DDPM denoising process of each step using a series of edit-friendly noise maps, and leverage DPM Solver++ to accelerate this process, enabling efficient sampling with semantic consistency. Compared to existing methods, our framework enables the efficient generation of adversarial examples that exhibit minimal discernible semantic changes. Consequently, we for the first time introduce Semantic-Consistent Adversarial Examples (SCAE). Extensive experiments and visualizations have demonstrated the high efficiency of SCA, particularly in being on average 12 times faster than the state-of-the-art attacks. Our code can be found at https://github.com/Pan-Zihao/SCA.
@@
Deep neural networks have excelled in learning features and achieving impressive performance across various tasks. However, they are vulnerable to small perturbations, known as adversarial samples, which raise significant security concerns for critical decision-making systems. To ensure human visual imperceptibility and maintain photorealism, adversarial examples must retain high semantic consistency with the original images, appearing natural and realistic despite minor perturbations [1, 2, 3, 4].
@@
A recent trend in adversarial attacks is the emergence of unrestricted adversarial attacks [5, 6, 2, 7], which alter semantic features such as color and texture to produce adversarial examples that are both visually plausible and highly effective [1, 8]. State-of-the-art methods [9, 10] utilize generative models like Stable Diffusion [11], leveraging their knowledge of natural image distributions to map images onto a latent space, where small perturbations are introduced to optimize the adversarial objective. While this approach allows for more flexible adversarial examples, it often causes noticeable semantic shifts, reducing stealthiness to human observers [12, 13]. A comparison of our method with previous work on semantic consistency is shown in Fig. 1. We argue that even small changes in the latent space can result in significant image alterations. Existing methods, constrained by norms, struggle to control these semantic variations. Furthermore, they typically require extensive time steps (e.g., 1000) for accurate image inversion, making the attack process both time-consuming and prone to uncontrolled semantic changes [9, 10].
@@
To tackle the challenges inherent in adversarial attacks, we introduce a novel approach called Semantic-Consistent Unrestricted Adversarial Attack (SCA), which leverages rich semantic information from the original images as a guide throughout the entire attack process. Our method begins by enhancing the inversion mechanism. Denoising Diffusion Probabilistic Models (DDPMs) typically produce images by progressively refining a series of white Gaussian noise samples. Departing from prior approaches [14, 15, 3] we propose an alternative latent noise space for DDPM, which allows for a wide range of editable operations through straightforward manipulation. This space enables the extraction of noise maps that are conducive to editing for a given image. In the context of generating unrestricted adversarial examples, an essential property of inversion is that perturbing the latent space while keeping the noise maps fixed should result in artifact-free images that retain a high degree of semantic consistency with the original image. To achieve this, the noise maps are designed to be statistically independent and to follow a standard normal distribution, akin to conventional sampling techniques. Our inversion process imprints the image more strongly onto the noise maps, thereby preserving semantics more effectively when traversing the low-dimensional manifold. This preservation is due to the fact that edit-friendly noise maps exhibit higher variances than their native counterparts. Additionally, Multimodal Large Language Models (MLLMs) have demonstrated a strong capacity for understanding a wide array of image semantics [16]. We leverage MLLMs to generate highly detailed image descriptions, which guide both the inversion and generation processes, thereby providing rich prior knowledge to maintain semantic consistency during perturbation in the latent space. Finally, we integrate DPM Solver++, an optimized numerical algorithm that accelerates the sampling process of DDPMs by reducing the number of required time steps to just 10–20. This significantly boosts the overall efficiency of our method.
@@ II Related Work
Adversarial attacks seek to maximize a model’s classification error while preserving image semantics. Below, we review key approaches.
@@
Norm-Bounded Attacks. Norm-bounded attacks are popular for their simplicity and effectiveness, using gradient-based optimization to maximize model loss. Techniques like FGSM [17], L-BFGS [18], PGD [19], and the CW attack [20] constrain perturbations within an norm, ensuring minimal modifications. However, these methods often introduce distortions detectable by humans and robust models, making them less effective in real-world applications. Norm constraints also limit flexibility, as norms do not always align with human perception, leading to visually suboptimal results.
Unrestricted Adversarial Attacks. Unrestricted attacks, which focus on natural transformations like shape, texture, and color, offer more flexibility than norm-bounded methods. Shape-based approaches [21, 22] alter image geometry but struggle to maintain realism. Texture-based attacks [5, 23] modify perceptual features but often produce unnatural results. Color-based attacks [1, 5, 24, 25, 26, 27] achieve photorealism but lack flexibility and transferability across models. These approaches balance realism and adversarial effectiveness, but each comes with trade-offs.
Latent Space-Based Attacks. Recent methods exploit generative models like GANs [13, 28, 29, 12] and diffusion models [30, 31, 9, 32] to manipulate the latent space and generate adversarial images. While offering control over image generation, these approaches often introduce noticeable semantic shifts, reducing stealthiness. Our framework addresses this by maintaining semantic consistency in adversarial examples, enhancing both stealth and practical applicability.
@@ V-B Semantic Consistency Comparison
In order to comprehensively evaluate the semantic consistency preservation ability of current unrestricted adversarial attack methods from multiple perspectives, we introduce four indicators, namely CLIP Score, SSIM, PSNR and MSE. The specific results are shown in Table I.
@@ V-C Comparison of Attack Success Rates
| Attack |
TReS | ||||||||
| Clean | 5.15 | 0.667 | 4.07 | 52.66 | 82.01 | ||||||||
| SAE | 5.05 | 0.597 | 3.79 | 47.24 | 71.88 | ||||||||
| ADef | 4.89 | 0.608 | 3.89 | 47.39 | 72.10 | ||||||||
| cAdv | 4.97 | 0.623 | 3.87 | 48.32 | 75.12 | ||||||||
| tAdv | 4.83 | 0.525 | 3.78 | 44.71 | 67.07 | ||||||||
| ACE | 5.12 | 0.648 | 3.96 | 50.49 | 77.25 | ||||||||
| ColorFool | 5.24 | 0.662 | 4.05 | 52.27 | 78.54 | ||||||||
| NCF | 4.96 | 0.634 | 3.87 | 50.33 | 74.10 | ||||||||
| AdvST | 5.12 | 0.679 | 3.99 | 56.33 | 79.21 | ||||||||
| ACA | 5.54 | 0.691 | 4.37 | 56.08 | 85.11 | ||||||||
| SCA(Ours) | 5.32 | 0.691 | 4.47 | 58.62 | 82.05 |
In this section, we conduct a detailed quantitative assessment of the image quality for adversarial examples, following the approach taken in prior works such as [25], [26], and [9]. Specifically, we employ non-reference perceptual image quality metrics to evaluate the visual realism and aesthetic appeal of the adversarial images generated by different methods. The metrics we use include NIMA, HyperIQA, MUSIQ, and TReS, which are widely recognized for their ability to assess image quality in the absence of a ground truth reference.
To provide a robust evaluation, we follow the PyIQA framework and apply models trained on large-scale aesthetic datasets. NIMA-AVA and MUSIQ-AVA are trained on the AVA dataset, while MUSIQ-KonIQ is trained on the KonIQ-10K dataset, which ensures a comprehensive evaluation of the generated adversarial examples. As shown in Table III, our method consistently achieves image quality scores that are comparable to ACA, outperforming other methods such as ColorFool and NCF. This superior performance is likely attributable to the fact that both SCA and ACA leverage Stable Diffusion, a state-of-the-art generative model known for its ability to produce highly realistic and visually appealing images.
@@ V-G Ablation Study
In the process of compute the gradients for optimization based on In the gradient optimization process using (10) and (11), we apply the random gradient-free (RGF) method. To evaluate its effectiveness, we conducted ablation tests under three conditions: without gradient estimation (”NE”), using the Skip Gradient method (”SG”), and using RGF. Table V shows the results. Our SCA performs poorly without gradient estimation and with the Skip Gradient method. The ASR for ”No estimation” is similar to the clean image, showing that we can’t ignore the diffusion model’s sampling and directly optimize with the classifier’s loss in the latent space. The Skip Gradient method also underperforms, likely due to the need for more sample steps, as we use complex optimizations to speed up sampling. Additionally, our new inversion method differs from prior approximations, validating the importance of the RGF method in our framework.

### v3

@@ Abstract
Deep neural network based systems deployed in sensitive environments are vulnerable to adversarial attacks. Unrestricted adversarial attacks typically manipulate the semantic content of an image (e.g., color or texture) to create adversarial examples that are both effective and photorealistic. Recent works have utilized the diffusion inversion process to map images into a latent space, where high-level semantics are manipulated by introducing perturbations. However, they often results in substantial semantic distortions in the denoised output and suffers from low efficiency. In this study, we propose a novel framework called Semantic-Consistent Unrestricted Adversarial Attacks (SCA), which employs an inversion method to extract edit-friendly noise maps and utilizes Multimodal Large Language Model (MLLM) to provide semantic guidance throughout the process. Under the condition of rich semantic information provided by MLLM, we perform the DDPM denoising process of each step using a series of edit-friendly noise maps, and leverage DPM Solver++ to accelerate this process, enabling efficient sampling with semantic consistency. Compared to existing methods, our framework enables the efficient generation of adversarial examples that exhibit minimal discernible semantic changes. Consequently, we for the first time introduce Semantic-Consistent Adversarial Examples (SCAE). Extensive experiments and visualizations have demonstrated the high efficiency of SCA, particularly in being on average 12 times faster than the state-of-the-art attacks. Our research can further draw attention to the security of multimedia information.
@@
Deep neural networks have excelled in learning features and achieving impressive performance across various tasks. However, they are vulnerable to small perturbations, known as adversarial samples, which raise significant security concerns for critical decision-making systems [1, 2, 3, 4, 5, 6, 7, 8]. With the explosive growth of digital information, they have triggered security risks such as face recognition [9, 10], voice recognition [11, 12, 13] and information forensics [14, 15].
@@
To ensure human visual imperceptibility and maintain photorealism, adversarial examples must retain high semantic consistency with the original images, appearing natural and realistic despite minor perturbations [16, 17, 18, 19]. A recent trend in adversarial attacks is the emergence of unrestricted adversarial attacks [20, 21, 17, 22], which alter semantic features such as color and texture to produce adversarial examples that are both visually plausible and highly effective [16, 23]. State-of-the-art methods [24, 25] utilize generative models like Stable Diffusion [26], leveraging their knowledge of natural image distributions to map images onto a latent space, where small perturbations are introduced to optimize the adversarial objective. While this approach allows for more flexible adversarial examples, it often causes noticeable semantic shifts, as shown in Fig. 2, reducing stealthiness to human observers [27, 28]. A comparison of our method with previous work on semantic consistency is shown in Fig. 1. We argue that even small changes in the latent space can result in significant image alterations. Existing methods, constrained by norms, struggle to control these semantic variations. Furthermore, they typically require extensive time steps (e.g., 1000) for accurate image inversion, making the attack process both time-consuming and prone to uncontrolled semantic changes [24, 25].
@@
To tackle the challenges inherent in adversarial attacks, we introduce a novel approach called Semantic-Consistent Unrestricted Adversarial Attack (SCA), which leverages rich semantic information from the original images as a guide throughout the entire attack process. Our method begins by enhancing the inversion mechanism. Denoising Diffusion Probabilistic Models (DDPMs) typically produce images by progressively refining a series of white Gaussian noise samples. Departing from prior approaches [29, 30, 18] we propose an alternative latent noise space for DDPM, which allows for a wide range of editable operations through straightforward manipulation. This space enables the extraction of noise maps that are conducive to editing for a given image. In the context of generating unrestricted adversarial examples, an essential property of inversion is that perturbing the latent space while keeping the noise maps fixed should result in artifact-free images that retain a high degree of semantic consistency with the original image. To achieve this, the noise maps are designed to be statistically independent and to follow a standard normal distribution, akin to conventional sampling techniques. Our inversion process imprints the image more strongly onto the noise maps, thereby preserving semantics more effectively when traversing the low-dimensional manifold. This preservation is due to the fact that edit-friendly noise maps exhibit higher variances than their native counterparts. Additionally, Multimodal Large Language Models (MLLMs) have demonstrated a strong capacity for understanding a wide array of image semantics [31]. We leverage MLLMs to generate highly detailed image descriptions, which guide both the inversion and generation processes, thereby providing rich prior knowledge to maintain semantic consistency during perturbation in the latent space. Finally, we integrate DPM Solver++ [32], an optimized numerical algorithm that accelerates the sampling process of DDPMs by reducing the number of required time steps to just 10–20. This significantly boosts the overall efficiency of our method.
@@ II Related Work
Adversarial attacks aim at maximizing the classification error of a target model without changing the semantics of the images. In this section, we review the types of attacks and existing works.
@@
Norm-Bounded Attacks. Norm-bounded attacks are popular for their simplicity and effectiveness, using gradient-based optimization to maximize model loss [1, 7, 6]. Techniques like FGSM [33], L-BFGS [34], PGD [35], and the CW attack [36] constrain perturbations within an norm, ensuring minimal modifications. However, these methods often introduce distortions detectable by humans and robust models, making them less effective in real-world applications. Norm constraints also limit flexibility, as norms do not always align with human perception, leading to visually suboptimal results.
Unrestricted Adversarial Attacks. As limitations of normbounded attacks become increasingly apparent, a growing interest in unrestricted adversarial attacks has emerged. These attacks aim to replace the traditional small norm perturbations with unrestricted but natural transformations, making them more applicable in real-world scenarios. Unrestricted adversarial attacks typically rely on semantic transformations, such as changes in shape, texture, or color, to craft adversarial examples. Shape-based unrestricted attacks, such as those by Alaifari et al. [37] and Xiao et al. [38], iteratively introduce minor deformations to an image using gradient descent to generate adversarial examples. These methods exploit subtle changes in the geometry of the image, but are often limited by the complexity of maintaining natural-looking outcomes. Following shape-based approaches, texture-based attacks emerged as an alternative. Bhattad et al. [20], Qiu et al. [39] and Mou et al. [2] proposed methods that manipulate an image’s texture or style to induce adversarial behavior. While texture-based attacks provide an effective means of altering perceptual features, they often lead to results that are visually unnatural, thereby diminishing their practical utility and adversarial transferability across different models. The unnatural distortions generated by these approaches highlight the challenge of balancing perceptual realism with adversarial effectiveness.
To address these challenges, researchers turned to colorbased unrestricted attacks, which aim to manipulate pixel values along specific dimensions to maintain a higher degree of realism. Works by Hosseini et al. [16], Bhattad et al. [20], Zhao et al. [40], Shamsabadi et al. [41], and Yuan et al. [42] focus on altering color properties to generate adversarial examples that appear more natural to human observers. Despite achieving photorealism, these attacks often sacrifice flexibility in generating diverse adversarial examples. The reliance on subjective visual intuitions, coupled with the use of objective metrics like perceptual similarity, constrains the range of possible adversarial transformations. Consequently, color-based attacks frequently exhibit limited transferability to other models, reducing their effectiveness in broader adversarial settings.
In summary, while shape-, texture-, and color-based unrestricted attacks have contributed significantly to the field of adversarial machine learning, each method comes with inherent trade-offs between realism, flexibility, and adversarial transferability. Our work builds on these insights, aiming to overcome the limitations of existing approaches by developing a more flexible and transferable framework for generating natural adversarial examples without compromising photorealism.
Latent Space-Based Attacks. Recent advancements in adversarial attack techniques have focused on manipulating the latent space of generative models to craft adversarial images. For instance, Zhao et al.[28], Lin et al. [43], Hu et al. [44], Lapid et al. [45], and Lau et al. [27] leveraged Generative Adversarial Networks (GANs), fine-tuning the generator to produce adversarial images. These methods typically adjust the generator’s parameters to subtly modify the image’s latent representation, thus creating visually plausible yet adversarial content. On the other hand, diffusion models have also been employed for similar purposes. Xue et al. [46], Wang et al. [47], Chen et al. [29], Liu et al. [48], and Chen et al. [24] focused on optimizing the U-Net structure or injecting learned noise into the latent space to generate adversarial examples. Although diffusion models offer enhanced flexibility in controlling image generation processes, a common issue across both GAN-based and diffusion-based approaches is the pronounced semantic shift in the adversarial examples. These shifts can often make the manipulations more perceptible to humans, thus undermining the stealthiness of the attack. In contrast, our proposed framework addresses these limitations by generating adversarial examples that maintain a higher degree of semantic consistency, ensuring the resulting images exhibit minimal perceptible alterations while still achieving the desired adversarial effect. This improvement in stealthiness enhances the practical viability of adversarial attacks in real-world applications.
@@ V-B Semantic Consistency Comparison
In order to comprehensively evaluate the semantic consistency preservation ability of current unrestricted adversarial attack methods from multiple perspectives, we introduce four indicators, namely CLIP Score, SSIM, PSNR and MSE. The specific results are shown in Table IV. PSNR (Peak Signal to Noise Ratio) is a metric for evaluating image quality. The unit of PSNR is dB. The larger the value, the less image distortion. However, many experimental results show that the PSNR score cannot be completely consistent with the visual quality seen by the human eye. It is possible that a higher PSNR image looks worse than a lower PSNR image. This is because the sensitivity of the human eye to errors is not absolute, and its perception results will vary due to many factors. In order to address the limitations of PSNR, we introduced LPIPS and SSIM. The range of SSIM values is [0,1], and the larger the value, the more similar the images are. If two images are exactly the same, the SSIM value is 1. LPIPS (Learned Perceptual Image Patch Similarity), also known as perceptual loss, is used to measure the difference between two images. This metric learns the reverse mapping from the generated image to the ground truth, forces the generator to learn the reverse mapping of reconstructing the real image from the fake image, and prioritizes the perceptual similarity between them. LPIPS is more in line with human perception than traditional methods. The lower the LPIPS value, the more similar the two images are, and vice versa, the greater the difference.
In addition, we also introduced CLIP Score as a more advanced semantic evaluation metric. As we all know, CLIP has been trained on more than 400 million image-text pairs and has strong understanding capabilities. So we measure semantic consistency by calculating the cosine similarity between the image embeddings processed by CLIP’s visual encoder.
@@ V-C Comparison of Attack Success Rates
| Attack |
TReS | ||||||||
| Clean | 5.15 | 0.667 | 4.07 | 52.66 | 82.01 | ||||||||
| SAE | 5.05 | 0.597 | 3.79 | 47.24 | 71.88 | ||||||||
| ADef | 4.89 | 0.608 | 3.89 | 47.39 | 72.10 | ||||||||
| cAdv | 4.97 | 0.623 | 3.87 | 48.32 | 75.12 | ||||||||
| tAdv | 4.83 | 0.525 | 3.78 | 44.71 | 67.07 | ||||||||
| ACE | 5.12 | 0.648 | 3.96 | 50.49 | 77.25 | ||||||||
| ColorFool | 5.24 | 0.662 | 4.05 | 52.27 | 78.54 | ||||||||
| NCF | 4.96 | 0.634 | 3.87 | 50.33 | 74.10 | ||||||||
| AdvST | 5.12 | 0.679 | 3.99 | 56.33 | 79.21 | ||||||||
| ACA | 5.54 | 0.691 | 4.37 | 56.08 | 85.11 | ||||||||
| SCA(Ours) | 5.32 | 0.691 | 4.47 | 58.62 | 82.05 |
@@
In this section, we conduct a detailed quantitative assessment of the image quality for adversarial examples, following the approach taken in prior works such as [41], [42], and [24]. Specifically, we employ non-reference perceptual image quality metrics to evaluate the visual realism and aesthetic appeal of the adversarial images generated by different methods. The metrics we use include NIMA [68], HyperIQA [69], MUSIQ [70], and TReS [71], which are widely recognized for their ability to assess image quality in the absence of a ground truth reference.
To provide a robust evaluation, we follow the PyIQA framework and apply models trained on large-scale aesthetic datasets. NIMA-AVA and MUSIQ-AVA are trained on the AVA [72] dataset, while MUSIQ-KonIQ is trained on the KonIQ-10K [73] dataset, which ensures a comprehensive evaluation of the generated adversarial examples. As shown in Table III, our method consistently achieves image quality scores that are comparable to ACA, outperforming other methods such as ColorFool and NCF. This superior performance is likely attributable to the fact that both SCA and ACA leverage Stable Diffusion, a state-of-the-art generative model known for its ability to produce highly realistic and visually appealing images.
@@ V-G Ablation Study
In the process of compute the gradients for optimization based on In the gradient optimization process using (13) and (15), we apply the random gradient-free (RGF) method. To evaluate its effectiveness, we conducted ablation tests under three conditions: without gradient estimation (NE), using the Skip Gradient method (SG), and using RGF. Table V shows the results. Our SCA performs poorly without gradient estimation and with the Skip Gradient method. The ASR for ”No estimation” is similar to the clean image, showing that we can’t ignore the diffusion model’s sampling and directly optimize with the classifier’s loss in the latent space. The Skip Gradient method also underperforms, likely due to the need for more sample steps, as we use complex optimizations to speed up sampling. Additionally, our new inversion method differs from prior approximations, validating the importance of the RGF method in our framework.